# Compiler Optimization via LLM Reasoning for Efficient Model Serving

### v1

@@ 6 Conclusion
The cost and complexity of compiling neural workloads remain major barriers to scalable and accessible model serving. Traditional compilers struggle with the vast, interdependent transformation space, while existing stochastic methods often lack sample efficiency and contextual awareness. We introduced Reasoning Compiler, a novel framework that formulates compiler optimization as a sequential, context-aware decision process. Using LLM as proposal generators and structured Monte Carlo Tree Search (MCTS) for guided exploration, Reasoning Compiler efficiently navigates the optimization space by combining LLM reasoning with performance feedback. Our results show that LLM-guided compilation, without any retraining, can substantially improve sample efficiency.
@@ Appendix A LLM Prompt Example
@tvm.script.ir_module class MyModule: @T.prim_func def main( A: T.Buffer((1, 16, 7168), "float32"), B: T.Buffer((7168, 2048), "float32"), C: T.Buffer((1, 16, 2048), "float32"), ): T.func_attr({"global_symbol": "main", "tir.noalias": True}) for b, t, j, k in T.grid(1, 16, 2048, 7168): with T.block("moe"): vb, vt, vj, vk = T.axis.remap("SSSR", [b, t, j, k]) with T.init(): C[vb, vt, vj] = T.float32(0) C[vb, vt, vj] += A[vb, vt, vk] * B[vk, vj]
@@
You are an code optimization assistant performing Monte Carlo Tree Search (MCTS) on a given code to improve performance. Each code has a corresponding history of transformations and predicted cost. You are given the code of the selected node and its immediate parent. Their main differences include: Loop shapes: Current: for b_1, t_1, j_1 in T.grid(1, 2, 8), and then j_2_init in T.grid(1), and later j_2 in T.grid(1). Index example: vj = ... + j_1 * 64 + j_2 * 64 Parent: for b_1, t_1, j_1 in T.grid(1, 2, 2), then j_2_init in T.grid(4), and later j_2 in T.grid(4). Index example: vj = ... + j_1 * 256 + j_2 * 64 Tile decisions: Current: sample_perfect_tile(..., decision=[4, 8, 1, 64]) Parent: sample_perfect_tile(..., decision=[4, 2, 4, 64]) Performance estimates: Current: 0.773 Parent: 0.313 Available transformations: TileSize, Parallel, ComputeLocation, Unroll Task Analyze the IR, trace, and predicted scores. Then propose a sequence of transformations (you may repeat any) to potentially improve performance. Output the final suggested transformations in a single line starting with "Transformations to apply:..." For example: Transformations to apply: TileSize, Unroll, ...
@@
Transformations to apply: TileSize, TileSize, Unroll, Parallel, TileSize
@@ Appendix B Sample Efficiency and Speedup Results
Table 1 presents the relative speedup of three methods—Evolutionary Search, MCTS, and LLM-Guided MCTS —evaluated across the different benchmarks. Speedup is measured as the ratio of execution time for the unoptimized code to that of the optimized code after applying a given number of transformation proposals. The table captures performance as a function of the number of samples explored. Higher values indicate more effective optimization. For instance, LLM-Guided MCTS consistently achieves higher speedups with fewer samples, demonstrating superior sample efficiency and faster convergence compared to MCTS and Evolutionary Search.
Llama3-8B (Self-Attention Layer) | Number of Samples | 18 | 36 | 72 | 144 | 192 | 600 | 900 | 1632 | 5952 |
@@
DeepSeek-R1 (MoE Layer) | Number of Samples | 36 | 54 | 72 | 144 | 192 | 600 | 900 | 1632 | 3000 |
@@
FLUX (Self-Attention Layer) | Number of Samples | 36 | 54 | 72 | 150 | 200 | 600 | 1000 | 1500 | 3000 |
@@
FLUX (Convolution) | Number of Samples | 36 | 54 | 72 | 150 | 200 | 400 | 600 | 800 | 1000 |
Evolutionary Search | 2.08 | 2.15 | 2.11 | 2.15 | 2.19 | 2.29 | 2.32 | 2.44 | 2.44 | |
MCTS | 1.96 | 2.13 | 2.13 | 2.18 | 2.37 | 2.38 | 2.38 | 2.39 | 2.40 | |
MCTS + LLM | 2.21 | 2.21 | 2.30 | 2.29 | 2.36 | 2.47 | 2.47 | 2.51 | 2.51 |
@@ Appendix C Impact of LLM Choice and Reasoning Strategy
As a continuation of Figure 5(a), Table 2 reports speedup over unoptimized code on three additional benchmarks: DeepSeek-R1 MoE Layer, FLUX Self-Attention Layer, and FLUX Convolution Layer. Each block of the table corresponds to a different benchmark and shows the best speedup achieved by LLM-Guided MCTS as a function of the number of schedules sampled using the reasoning model listed in the table. Rows compare different reasoning models used for API call generation, including both proprietary (e.g., GPT-4o mini, OpenAI o1-mini) and open-source models (e.g., Llama3.3-Instruct, DeepSeek-Distill). Across all benchmarks, the results show that more capable models—those that are larger or instruction-tuned—consistently achieve higher speedups with fewer samples. For example, Llama3.3-Instruct (70B) and DeepSeek-Distill (32B) achieve near-maximal speedup within the first 72–150 samples, while smaller models such as DeepSeek-Distill (7B) or Llama3.1-Instruct (8B) reach similar performance more gradually. These results validate the generality of our findings: the use of context-aware LLMs accelerates convergence in LLM-Guided MCTS across diverse code domains. Moreover, the performance of open-source models is competitive with proprietary alternatives, further supporting the accessibility and reproducibility of our method.
Llama3-8B (Self-Attention Layer) | Number of Samples | 18 | 36 | 72 | 150 | 200 | 600 |
@@
DeepSeek-R1 (MoE Layer) | Number of Samples | 18 | 36 | 72 | 150 | 200 | 600 |
@@
FLUX (Self-Attention Layer) | Number of Samples | 18 | 36 | 72 | 150 | 200 | 600 |
@@
FLUX (Convolution) | Number of Samples | 18 | 36 | 72 | 150 | 200 | 600 |
@@ Appendix D Impact of Historical Trace Depth on Optimization Efficiency
Results show that increasing the historical context generally improves sample efficiency across all benchmarks. For example, on DeepSeek-R1 MoE Layer, adding one more ancestral node boosts early performance significantly, achieving a 9.39 speedup at just 18 samples compared to 6.14 for the shallower context. Similarly, on Llama3-8B Self-Attention Layer, the extended context leads to a higher final speedup (11.87 vs. 11.33) and earlier convergence. The performance gains, while smaller, are also consistent on FLUX Self-Attention Layer and FLUX Convolution Layer, with improvements observed across all sample budgets. These findings confirm that providing richer historical context enables the LLM to make more informed decisions at each step of the search, ultimately enhancing the sample efficiency of LLM-Guided MCTS.
Llama3-8B (Self-Attention Layer) | Number of Samples | 18 | 36 | 72 | 150 | 200 | 600 |
@@
DeepSeek-R1 (MoE Layer) | Number of Samples | 18 | 36 | 72 | 150 | 200 | 600 |
@@
FLUX (Self-Attention Layer) | Number of Samples | 18 | 36 | 72 | 150 | 200 | 600 |
@@
FLUX (Convolution) | Number of Samples | 18 | 36 | 72 | 150 | 200 | 600 |
@@ Appendix E Ablations of MCTS Branching Factor
To determine the value of MCTS branching factor (), we ablate on and . In Table 4, results show that when branching factor , LLM-Guided MCTS is more sample-efficient than when . Our choice of aligns with prior works [14, 15]. If a higher branching factor is chosen, then there are more possible next steps, which require more sampling effort (i.e., more simulations) to cover these expanded possibilities at the same level of thoroughness.
Llama3-8B (Self-Attention Layer) | Number of Samples | 18 | 36 | 72 | 150 | 200 | 600 |
@@
DeepSeek-R1 (MoE Layer) | Number of Samples | 18 | 36 | 72 | 150 | 200 | 600 |
@@
FLUX (Self-Attention Layer) | Number of Samples | 18 | 36 | 72 | 150 | 200 | 600 |
@@
FLUX (Convolution) | Number of Samples | 18 | 36 | 72 | 150 | 200 | 600 |

### v3

@@ 6 Conclusion
Compiling neural workloads remains a bottleneck for scalable model serving: traditional compilers struggle with combinatorial transformation spaces, and the state-of-the-art neural compilers rely on stochastic search, lacking sample efficiency and contextual awareness. This paper introduced Reasoning Compiler, a novel framework that formulates compiler optimization as a sequential, context-aware decision process, pairing LLM‑generated proposals with MCTS and performance feedback to reason and navigate through the optimization space efficiently. By enabling LLM reasoning in the compiler optimization process, we achieve a leap from randomized search to informed and guided compilation. Our results show that Reasoning Compiler consistently yields faster runtimes with markedly fewer evaluations without any retraining. These gains directly translate to reduced operational cost of LLM services, lower energy usage per query, improved system responsiveness, more agile model deployment, faster model training, and accelerated innovation cycles, among other benefits. Looking ahead, the same LLM that guides compilation can accelerate its own inferencing, creating a virtuous, self‑optimizing cycle in which sped-up LLMs enable more efficient transformations and progressively better models and services.
Acknowledgments
We thank the anonymous reviewers for their insightful feedback. This work was in part supported by the National Science Foundation (NSF) award CCF #2107598. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes not withstanding any copyright notation thereon. The views contained herein are those of the authors and should not be interpreted as necessarily representing the official policies or endorsements, either expressed or implied by the U.S. Government.
@@ Appendix A LLM Prompt Example
@tvm.script.ir_module
class MyModule:
@T.prim_func
def main(
A: T.Buffer((1, 16, 7168), "float32"),
B: T.Buffer((7168, 2048), "float32"),
C: T.Buffer((1, 16, 2048), "float32"),
T.func_attr({"global_symbol": "main", "tir.noalias": True})
for b, t, j, k in T.grid(1, 16, 2048, 7168):
with T.block("moe"):
vb, vt, vj, vk = T.axis.remap("SSSR", [b, t, j, k])
with T.init():
C[vb, vt, vj] = T.float32(0)
C[vb, vt, vj] += A[vb, vt, vk] * B[vk, vj]
@@
You are a code optimization assistant performing Monte Carlo Tree Search (MCTS) on a given code to improve performance. Each code has a corresponding history of transformations and predicted cost. You are given the code of the selected node and its immediate parent. Their main differences include: Loop shapes: Current: for b_1, t_1, j_1 in T.grid(1, 2, 8), and then j_2_init in T.grid(1), and later j_2 in T.grid(1). Index example: vj = ... + j_1 * 64 + j_2 * 64 Parent: for b_1, t_1, j_1 in T.grid(1, 2, 2), then j_2_init in T.grid(4), and later j_2 in T.grid(4). Index example: vj = ... + j_1 * 256 + j_2 * 64 Tile decisions: Current: sample_perfect_tile(..., decision=[4, 8, 1, 64]) Parent: sample_perfect_tile(..., decision=[4, 2, 4, 64]) Performance estimates: Current: 0.773 Parent: 0.313 Available transformations: TileSize, Parallel, ComputeLocation, Unroll Task Analyze the IR, trace, and predicted scores. Then propose a sequence of transformations (you may repeat any) to potentially improve performance. Output your reasoning and your suggested transformations. For example, your answer should be in the following format: Reasoning: This code still has large loop extents, so I’d tile it twice differently, then unroll. Transformations to apply: TileSize, TileSize, Unroll.
@@
Reasoning: The current schedule tiles the j-axis as 2048=4×8×1×64, whereas the parent uses 4×2×4×64. The higher predicted score of the current node (0.773 vs. 0.313) suggests that expanding outer parallel granularity (j_1=8) helped, but collapsing j_2 to 1 removed a small register-level tile, limiting unrolling and hurting reuse within each j_3=64 chunk. I would (1) reintroduce a non-unit j_2 while shrinking j_1 so that j_1×j_2˜=8; for instance, [4,4,2,64] keeps enough outer tasks but creates a j_2×j_3=128-wide microtile that unrolls cleanly. (2) Tile the 7168-step k-reduction to align with j_3=64, e.g., 7168 = 7×16×64; the 16-step inner slice can be unrolled to cut loop overhead and improve register reuse across the j_3 tile. (3) Move the compute of block "moe" to the j_1 level so each j_1 tile initializes/accumulates C locally, shortening live ranges and reducing write-back pressure. (4) Parallelize the outer j_1 (or t_1 if j_1 becomes too small) to utilize cores without touching the reduction. (5) Unroll the small j_2 loop and the k-inner microtile. These changes retain the beneficial outer structure of the current node while restoring a compact inner tile for ILP and cache locality, which should further improve throughput. Transformations to apply: TileSize, TileSize, ComputeLocation, Parallel, Unroll, Unroll.
@@ Appendix B Sample Efficiency and Speedup Results
Table 3 presents the relative speedup of three methods—Evolutionary Search, MCTS, and LLM-Guided MCTS—evaluated across the different benchmarks. Speedup is measured as the ratio of execution time for the unoptimized code to that of the optimized code after applying a given number of transformation proposals. The table captures performance as a function of the number of samples explored. Higher values indicate more effective optimization. For instance, LLM-Guided MCTS consistently achieves higher speedups with fewer samples, demonstrating superior sample efficiency and faster convergence compared to MCTS and Evolutionary Search. This table corresponds to Figure 3 in the paper.
| Llama-3-8B Attention Layer | Number of Samples | 18 | 36 | 72 | 144 | 192 | 600 | 900 | 1632 | 5952 |
@@
| DeepSeek-R1 MoE Layer | Number of Samples | 36 | 54 | 72 | 144 | 192 | 600 | 900 | 1632 | 3000 |
@@
| FLUX Attention Layer | Number of Samples | 36 | 54 | 72 | 150 | 200 | 600 | 1000 | 1500 | 3000 |
@@
| FLUX Convolution Layer | Number of Samples | 36 | 72 | 150 | 200 | 400 | 600 | 1000 | 1600 | 3000 |
| Evolutionary Search | 2.08 | 2.11 | 2.15 | 2.19 | 2.29 | 2.32 | 2.44 | 2.45 | 2.55 | |
| MCTS | 2.11 | 2.13 | 2.18 | 2.37 | 2.38 | 2.38 | 2.44 | 2.44 | 2.45 | |
| MCTS + LLM | 2.21 | 2.30 | 2.29 | 2.36 | 2.47 | 2.47 | 2.51 | 2.55 | 2.58 | |
| Llama-4-Scout MLP Layer | Number of Samples | 20 | 50 | 100 | 250 | 400 | 600 | 1000 | 1500 | 3000 |
| Evolutionary Search | 1.36 | 2.28 | 3.61 | 5.59 | 5.59 | 5.75 | 5.76 | 5.94 | 5.94 | |
| MCTS | 1.76 | 2.51 | 4.05 | 5.41 | 7.83 | 8.13 | 8.58 | 8.90 | 8.90 | |
| MCTS + LLM | 12.74 | 12.74 | 12.74 | 12.75 | 12.75 | 13.24 | 13.26 | 13.52 | 13.79 |
@@ Appendix C Impact of LLM Choice and Reasoning Strategy
As a continuation of Figure 4(a), Table 4 reports speedup over unoptimized code on three additional benchmarks: DeepSeek-R1 MoE Layer, FLUX Attention Layer, and FLUX Convolution Layer. Each block of the table corresponds to a different benchmark and shows the best speedup achieved by LLM-Guided MCTS as a function of the number of schedules sampled using the reasoning model listed in the table. Rows compare different reasoning models used for API call generation, including both proprietary (e.g., GPT-4o mini, OpenAI o1-mini) and open-source models (e.g., Llama3.3-Instruct, DeepSeek-Distill). Across all benchmarks, the results show that more capable models—those that are larger or instruction-tuned—consistently achieve higher speedups with fewer samples. For example, Llama3.3-Instruct (70B) and DeepSeek-Distill (32B) achieve near-maximal speedup within the first 72–150 samples, while smaller models such as DeepSeek-Distill (7B) or Llama3.1-Instruct (8B) reach similar performance more gradually. These results validate the generality of our findings: the use of context-aware LLMs accelerates convergence in LLM-Guided MCTS across diverse code domains. Moreover, the performance of open-source models is competitive with proprietary alternatives, further supporting the accessibility and reproducibility of our method.
| Llama3-8B Attention Layer | Number of Samples | 18 | 36 | 72 | 150 | 200 | 600 |
@@
| DeepSeek-R1 MoE Layer | Number of Samples | 18 | 36 | 72 | 150 | 200 | 600 |
@@
| FLUX Attention Layer | Number of Samples | 18 | 36 | 72 | 150 | 200 | 600 |
@@
| FLUX Convolution Layer | Number of Samples | 18 | 36 | 72 | 150 | 200 | 600 |
@@ Appendix D Impact of Historical Trace Depth on Optimization Efficiency
Results show that increasing the historical context generally improves sample efficiency across all benchmarks. For example, on DeepSeek-R1 MoE Layer, adding one more ancestral node boosts early performance significantly, achieving a 9.39 speedup at just 18 samples compared to 6.14 for the shallower context. Similarly, on Llama-3-8B Attention Layer, the extended context leads to a higher final speedup (11.87 vs. 11.33) and earlier convergence. The performance gains, while smaller, are also consistent on FLUX Attention Layer and FLUX Convolution Layer, with improvements observed across all sample budgets. These findings confirm that providing richer historical context enables the LLM to make more informed decisions at each step of the search, ultimately enhancing the sample efficiency of LLM-Guided MCTS.
| Llama-3-8B Attention Layer | Number of Samples | 18 | 36 | 72 | 150 | 200 | 600 |
@@
| DeepSeek-R1 MoE Layer | Number of Samples | 18 | 36 | 72 | 150 | 200 | 600 |
@@
| FLUX Attention Layer | Number of Samples | 18 | 36 | 72 | 150 | 200 | 600 |
@@
| FLUX Convolution Layer | Number of Samples | 18 | 36 | 72 | 150 | 200 | 600 |
@@ Appendix E Ablations of MCTS Branching Factor
To determine the value of MCTS branching factor (), we ablate on and . In Table 6, results show that when branching factor , LLM-Guided MCTS is more sample-efficient than when . Our choice of aligns with prior works [branching1:cg:2007, ucb:ml:2002]. If a higher branching factor is chosen, then there are more possible next steps, which require more sampling effort (i.e., more simulations) to cover these expanded possibilities at the same level of thoroughness.
| Llama-3-8B Attention Layer | Number of Samples | 18 | 36 | 72 | 150 | 200 | 600 |
@@
| DeepSeek-R1 MoE Layer | Number of Samples | 18 | 36 | 72 | 150 | 200 | 600 |
@@
| FLUX Attention Layer | Number of Samples | 18 | 36 | 72 | 150 | 200 | 600 |
@@
| FLUX Convolution Layer | Number of Samples | 18 | 36 | 72 | 150 | 200 | 600 |